OpenStudy (anonymous):

can someone help me undersntad this example in yahoo:

OpenStudy (anonymous):

how did they choose 5 and 2?

OpenStudy (anonymous):

i am talking about the top answer on there

OpenStudy (anonymous):

using bahaskar-bouckner algorithem

OpenStudy (anonymous):

5/2 is an approximation of sqrt(5)

OpenStudy (anonymous):

so for this method we have to use 5/2 as a starting point

OpenStudy (anonymous):

no.

OpenStudy (anonymous):

well i mean in this case cause n=5 right?

OpenStudy (anonymous):

but if i wanted to find square root of 45, using the bhaskara method i would start with 45/2

OpenStudy (anonymous):

no try it with 6/2 :)

OpenStudy (anonymous):

your smart guy ehh:)

OpenStudy (anonymous):

just takes longer to get closer to the value if you start from farther away

OpenStudy (anonymous):

okay i got it broski, was just not sure about how to start it, cause am writing a program for it, but i like to actually work out the first few terms on paper, so i know what i am doing

OpenStudy (anonymous):

sqrt(45) is 3sqrt(5) so you'd probably start with 3*5/2

OpenStudy (anonymous):

I don't think it matters, I don't think you can diverge from the value if you pick a wild starting approximation... it just takes more steps to get a useful approximation..

OpenStudy (anonymous):

cool

OpenStudy (cwrw238):

theres an old method I'm which you split the number in pairs working from right as follows

OpenStudy (anonymous):

Brounker:=proc(n,m) local a,b,i; a[1]:=1; b[1]:=1; for i from 1 to m-1 do a[i+1]:=a[i]+b[i]*n; b[i+1]:=a[i]+b[i]; end do; return a[m]/b[m],evalf(abs(a[m]/b[m]-sqrt(n))); end proc;

OpenStudy (anonymous):

Basically this is the directions: For a positive integer n, an iteration is known as the Bhaskara-Brouncker algorithm for approximating √ n by rational numbers ai/ bi , where ai, and bi for i = 1, 2, . . . are integers defined by the recurrence formula ai+1 = ai + bi · n, bi+1 = ai + bi for i = 1, 2, . . . where a1 = b1 = 1. Write a program that, for input positive integers n and m, carries out the iteration for m steps and outputs am /bm and the error am bm − √ n

OpenStudy (cwrw238):